# Split Two-Periodic Aztec Diamond

### v2

@@ 4. Local Asymptotics of the Model
The goal of this section is to prove the local asymptotic results stated in Propositions 2.1 and 2.2. We begin in Section 4.1 by stating some necessary results about the behavior of the saddle functions. In particular, we need precise statements about the location of saddle points and the contours of steepest ascent and descent in the smooth and rough regions of the model. Once we have the necessary preliminary information, we prove Propositions 2.1 and 2.2 in Sections 4.2 and 4.3, respectively.
@@ 4.2.1. Asymptotics on the $\alpha$-side when $\beta<\alpha$
We start with the definition of $[I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)]_{i,j=0}^{1}$ presented in equation 27. Next, deform the $w$-contour so it will ultimately surround the branch cuts, which are depicted in Figure 12. We do so by pushing the $w$-contour to infinity, and then flipping it around so it surrounds the negative real axis. We shrink the contour to just around the branch cuts. Where there is not a branch cut along the negative real axis, the contours cancel out. We are left with two separate contours, $\gamma_{w,1}$ and $\gamma_{w,2}$. We also deform the $z$-contour to the steepest ascent contour through the saddle point $z_{2}^{*}$. We refer to this contour as $\gamma_{s,2}$ and it is described in Lemma 4.3. The resulting contours are depicted in Figure 13.
@@
In deforming the contours we should be aware of the residue at $z=w$ that results from crossing the contours, however this term drop out because of the matrix product that appears.
$F_{\alpha,1}(z)F_{\alpha,2}(z)=0$
Thus we are left with the equation,
@@ 4.3. Asymptotics of $I^{\alpha}_{2,1}$
We will now work towards the proof of Proposition 2.1. We begin by manipulating the double contour integral in $I^{\alpha}_{2,1}$ into a form that is more amenable to asymptotic analysis. We recall the form of $I^{\alpha}_{2,1}$, which we don’t write in terms of the saddle functions presently.
(83) $\Big{[}I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\\ \times r_{\alpha,1}(z)^{m-\frac{N}{2}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)$
We will repeat a procedure seen in the proof of Corollary 2.1 in Section 3.4. Namely, we will deform the $w$-contours through the branch cuts and onto the other sheet of the Riemann surface. In doing so, we gain a single contour integral from the residue at $z=w$.
(84) $\Big{[}I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}\\ \times r_{\alpha,1}(z)^{m-\frac{N}{2}}\frac{2F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m+m^{\prime})}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}$
This leads us to state and prove the following lemma,
Lemma 4.6.
For $m,\,m^{\prime},\,\xi,$ and $\xi^{\prime}$ defined by (23) and (24),
(85) $\frac{1}{2\pi}\Bigg{|}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m+m^{\prime})}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Bigg{|}\leq Ce^{-cN}$
Where $C$ is some constant matrix and $c$ is a positive constant.
Proof.
Using the definitions of $m$, $m^{\prime}$, $\xi$, and $\xi^{\prime}$ from equations (23) and (24) we have,
$m+m^{\prime}=2xN+x_{1}+x_{2}<N$
$\xi^{\prime}-\xi=y_{2}-y_{1}$
The single contour integral in equation (84) has no pole at $z=1$, since $N-(m+m^{\prime})>0$, so we can deform the contour to a circle of radius one centered at the origin,
(86) $\frac{1}{2\pi\textrm{i}}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m+m^{\prime})}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}$
Along this contour, $|r_{\alpha,2}(z)|<1$ and the matrix-valued function
$\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}$
is bounded and analytic, so we can bound the integral accordingly,
(87) $\frac{1}{2\pi}\Bigg{|}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m+m^{\prime})}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Bigg{|}\leq C|r_{\alpha,2}^{\text{max}}|^{N-(m+m^{\prime})}$
where $r_{\alpha,2}^{\text{max}}$ denotes the maximum value of $r_{\alpha,2}(z)$ on the unit circle and $C$ is some constant matrix. ∎
To understand the asymptotics of $I_{2,1}^{\alpha}$, we are left considering the asymptotics of the remaining double contour in equation (84). The asymptotics of this double contour integral does depend on the region of the model, so we will consider the asymptotics in the rough and smooth regions separately.
@@ 4.3.1. Asymptotics of $I^{\alpha}_{2,1}$ in the Rough Region
Here we will prove the asymptotics of the double contour integral in equation (84) for $(x,y)\in\mathfrak{R}$ and $x<\nicefrac{{1}}{{2}}$. This result, along with Lemma 4.6, will prove part (1) of Proposition 2.2.
@@
Where $\Gamma_{r}$ and $\Gamma_{s}$ are lines through the origin whose angles depend on the angle of the contours of steepest descent and ascent, respectively. Combining the results from equations 89, 93 and Lemma 4.6 we obtain the results presented in part (1) of Proposition 2.2. ∎

### v3

@@ 4. Local Asymptotics of the Model
The goal of this section is to prove the local asymptotic results stated in Propositions 2.2 and 2.3. We begin in Section 4.1 by stating some necessary results about the behavior of the saddle functions. In particular, we need precise statements about the location of saddle points and the contours of steepest ascent and descent in the smooth and rough regions of the model. Once we have the necessary preliminary information, we prove Propositions 2.1-2.3 in Sections 4.2, 4.3, and 4.4.
@@ 4.1. Preliminaries of Saddle Functions
4.2. Proof of Proposition 2.1
As stated below equations (24) and (25), we will assume the coordinates $(m,\xi)$ and $(m^{\prime},\xi^{\prime})$ are on the same side of the interface. Regardless of which side they are on, $J$ has the form,
(70) $\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{|N-(m^{\prime}+m)|}M_{\varepsilon}(z)$
Since we are assuming $x\neq\nicefrac{{1}}{{2}}$ it follows that $|N-(m+m^{\prime})|>0$, so we can deform the contour to a circle of radius one centered at the origin,
(71) $\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{|N-(m^{\prime}+m)|}M_{\varepsilon}(z)$
Along this contour, $|r_{\varepsilon,2}(z)|<1$ and the matrix-valued function $M_{\varepsilon}$ is bounded and analytic, so we can bound the integral accordingly,
(72) $\frac{1}{2\pi}\Bigg{|}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{N-(m+m^{\prime})}M_{\varepsilon}(z)\Bigg{|}\leq C|r_{\alpha,2}^{\text{max}}|^{N-(m+m^{\prime})}$
where $r_{\varepsilon,2}^{\text{max}}$ denotes the maximum value of $r_{\varepsilon,2}(z)$ on the unit circle and $C$ is some constant matrix. The result in Proposition 2.1 follows.
@@ 4.3.1. Asymptotics on the $\alpha$-side when $\beta<\alpha$
We start with the definition of $[I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)]_{i,j=0}^{1}$ presented in equation 28. Since the $w$-contours transverse the branch cuts, we can consider them as two separate contours, which we call $\gamma_{w,1}$ and $\gamma_{w,2}$. We also deform the $z$-contour to the steepest ascent contour through the saddle point $z_{2}^{*}$. We refer to this contour as $\gamma_{s,2}$ and it is described in Lemma 4.3. The resulting contours are depicted in Figure 13.
@@
We write,
@@ 4.4. Asymptotics of $I^{\alpha}_{2,1}$
The asymptotics of this double contour integral does depend on the region of the model, so we will consider the asymptotics in the rough and smooth regions separately. Before we do so, we will repeat a procedure seen in the proof of Corollary 2.1 in Section 3.4 on the double contour integral $I^{\alpha}_{2,1}$. Namely, we will deform the $w$-contours through the branch cuts and onto the other sheet of the Riemann surface. Doing so we obtain,
(87) $\Big{[}I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}\\ \times r_{\alpha,1}(z)^{m-\frac{N}{2}}\frac{2F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(z)$
@@ 4.4.1. Asymptotics of $I^{\alpha}_{2,1}$ in the Rough Region
Here we will assume $(x,y)\in\mathfrak{R}$ and $x<\nicefrac{{1}}{{2}}$ and prove part (1) of Proposition 2.3.
@@
Where $\Gamma_{r}$ and $\Gamma_{s}$ are lines through the origin whose angles depend on the angle of the contours of steepest descent and ascent, respectively. Combining the results from equations 89, 93 and Lemma LABEL:lem:decaypart we obtain the results presented in part (1) of Proposition 2.3. ∎
@@ Appendix A Explicit Statement of Coefficients
Data Availability. No datasets were generated or analyzed for this research.
Conflict of Interest. There are no relevant financial or non-financial interest to disclose.